OpenStudy (kelsey_luh_ya_for_life_yeah_me):

x-2/7=3

OpenStudy (kelsey_luh_ya_for_life_yeah_me):

help

OpenStudy (kelsey_luh_ya_for_life_yeah_me):

anyone

OpenStudy (anonymous):

what you need

OpenStudy (kelsey_luh_ya_for_life_yeah_me):

look at the question

OpenStudy (anonymous):

you answered it

OpenStudy (kelsey_luh_ya_for_life_yeah_me):

no

OpenStudy (kelsey_luh_ya_for_life_yeah_me):

thats what i need help on

OpenStudy (notarobot):

x-2/7=3 21=x-2 x=23